OpenStudy (anonymous):

I need help with this simultaneous equation y = x^2 + 1 2y + 2 = x^2 + 6x Anyone can show me, the steps? Thanks

OpenStudy (anonymous):

first substitute: 2(x^2 + 1) +2= X^2 +6x 2x^2 + 2 + 2 = x^2 + 6x 2x^2 + 4 = x^2 + 6x x^2 -6x + 4 = 0 then factor... which i dont feel like doing right now.

OpenStudy (anonymous):

y = x^2 + 1 y + 2 = x^2 + 6x subtract 2 from both sides y = x^2 + 6x -2 substitute for y in first equation x^2 + 6x -2= x^2 + 1 subtrasct x^2 from both sides 6x -2= 1 add 2 to both sides 6x = 3 divide by 6 on both sides x = 1/2

OpenStudy (anonymous):

napervillian: wrong problem again.

OpenStudy (anonymous):

y = 5/4

OpenStudy (anonymous):

be clear x = 1/2 y = 5/4

OpenStudy (anonymous):

Thanks gianfranco once again !!!, understand it now the process you did it and also thanks elizated :)

OpenStudy (anonymous):

Glad to be of help.

OpenStudy (anonymous):

One question the second line it says y + 2 but it was 2y

OpenStudy (anonymous):

Yea, sorry I was too lazy to finish but Gianfraco saved the day!

OpenStudy (anonymous):

Serrico, then my answer is wrong.

OpenStudy (anonymous):

y = x^2 + 1 2y = x^2 + 6x multiply equation 1 by 2 2y = 2x^2 + 2 equate right hand sides x^2 + 6x = 2x^2 + 2 subtract x^2 from both sides 6x = x^2 + 2 subtract 6x from both sides 0 = x^2 -6x + 2 complete the square x = 3 -sqrt7 [and y = 3]or x = 3 + sqrt 7[and y = 17]

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

Thank you again, for taking your time in resolving it again even thought you didn't have too but it makes sense now. Using the square root otherwise I get weird X

OpenStudy (anonymous):

its 2y + 2 not only 2y in "2y = x^2 + 6x "

OpenStudy (anonymous):

y=x^2 +1 2y+2=x^2+6x then multiply the first equation by 2, we got 2y=2x^2+2 2y+2=x^2+6x then substitute 2y in the first equation into the second one (2x^2+2) + 2 = x^2+6x 2x^2+4 = x^2+6 (subtract x^2 from both sides) x^2+4=6x x^2-6x+4=0 (Elizated was right) then solve x by yourself, M>A>R>C

OpenStudy (anonymous):

ELIZATED FTW